OpenStudy (dyna-marie99):

Which function has an inverse that is a function? b(x) = x2 + 3 d(x) = –9 m(x) = –7x p(x) = |x|

satellite73 (satellite73):

the one that is "one to one" in this case, the line

satellite73 (satellite73):

do you know which one is a non horizontal line?

OpenStudy (dyna-marie99):

No

satellite73 (satellite73):

dang do you know what "one to one" means?

satellite73 (satellite73):

"no" is a fine answer, just asking is all

OpenStudy (dyna-marie99):

Each input has its own output

satellite73 (satellite73):

sorta each input gets an output for every fuction one to one means each input gets an output no other input gets so for example \[p(x)=|x|\] is NOT one to one since for exaple \[p(-5)=p(5)\]

satellite73 (satellite73):

neither is \(x^2+3\) because \((-2)^2+3=2^2+3\)

OpenStudy (dyna-marie99):

Ok

satellite73 (satellite73):

so that leaves only two the constant function and the line with slope \(-7\)

satellite73 (satellite73):

pick the line with slope \(-7\) it is one to one, the others are not

OpenStudy (dyna-marie99):

Thank you

satellite73 (satellite73):

yw now i have a question do you ride a harley?

satellite73 (satellite73):

"dyna" just had me wondering

OpenStudy (dyna-marie99):

No, i don't that is my real name

satellite73 (satellite73):

oh ok cool just wondering if it had another meaning like satllite73 is my car

OpenStudy (dyna-marie99):

ohk